# Beyond Subspace Isolation: Many-to-Many Transformer for Light Field Image Super-resolution

### v1

@@ IV-G Ablation Study
In this section, we undertake a few ablation studies to understand the characteristic of M2MT-Net and its individual components. Note that the studies are carried out at the most challenging 4×4$\times$4 × scale using the EPFL dataset as it has the most samples.
IV-G1 Spatial and Angular Components
To evaluate our proposed M2MT’s role in the spatial subspace, we substitute it with a vanilla Transformer or a convolution and trained the network. The modified networks have similar sizes with the original M2MT-Net to ensure a fair comparison. As indicated in Table III, substituting the M2MT with a vanilla Transformer deteriorates the performance by 0.51 dB to 29.29 dB. Opting for a convolution results in a further decline, with a drop of 0.78 dB to 29.02 dB. These outcomes affirm M2MT’s efficacy as a feature extractor in the spatial subspace than other alternatives when paired with its angular subspace counterpart.
We also train a model replacing the angular Transformer with a convolution. Surprisingly, this variant achieves 29.42 dB PSNR, slightly outperforming Transformer-based competitors like LFT and EPIT. This underscores the robust adaptability of the M2MT, even when paired with less potent components.
IV-G2 Number of Correlation Blocks
We explore the performance of M2MT-Net with varying the number of correlation blocks $n_{2}$. Table IV shows that M2MT-Net peaks in performance at $n_{2}=8$. Reducing the number leads to a gradual performance decline, and increasing it doesn’t guarantee performance improvement due to the potential risk of overfitting.
TABLE III: Ablation study on altering components in correlation blocks. The best PSNR/SSIM are in bold.
@@ IV-G2 Number of Correlation Blocks
TABLE IV: Ablation study on varing the number of correlation blocks. The best PSNR/SSIM are in bold.
Number 6 7 8 9 10
PSNR 29.60 29.69 29.80 29.65 29.66
SSIM 0.9257 0.9270 0.9277 0.9270 0.9271
V Conclusion
@@
In this paper, we have revealed the prevalent challenge of subspace isolation caused by the One-to-One scheme and present the novel concept of Many-to-Many Transformers (M2MT) as a new scheme to address this issue. The proposed M2MT is empowered with complete access to all pixels across all SAIs in a LF image to capture comprehensive long-range correlation dependencies. With M2MT as a pivotal component, we have proposed a simple yet effective M2MT-Net for LFSR. Extensive experiments on various public datasets have demonstrated that M2MT-Net surpasses state-of-the-art methods to a considerable extent. Its superiority is evidenced by visual interpretability in our in-depth analysis using the LAM technique, which highlights that M2MT involves a substantially broader range of pixels across wider SAIs beyond subspace isolation, signifying its truly global context and a more comprehensive modeling of correlation dependencies.
Looking ahead, there are some promising directions for improving M2MT in future works. Firstly, extending M2MT to EPI subspaces could potentially reinforce M2MT’s capacity to process LF images with large disparities like the STFgantry dataset [64], akin to EPIT [32]. However, the unique characteristics of EPIs present distinct challenges that necessitate further investigation. Secondly, M2MT-Net currently carries a relatively substantial model size when compared to other Transformer-based methods, despite its commendable speed. This is primarily attributed to the correlation encoding and decoding processes within M2MT. It would be advantageous to explore lightweight alternatives to achieve a balanced trade-off between computational efficiency and performance in future research endeavors.

### v2

@@ IV-G Ablation Study
TABLE III: Ablation study on altering components in correlation blocks. The best and second-best PSNR/SSIM are in bold and underlined.
@@
In this section, we undertake a few ablation studies to understand the characteristics of M2MT-Net and its individual components. Note that the studies are carried out at the most challenging 4×4$\times$4 × scale using the EPFL dataset with the most samples.
IV-G1 Spatial and Angular Components
To evaluate our proposed M2MT’s role in the spatial subspace, we substitute it with a vanilla Transformer or a convolution and train the network. The modified networks have similar sizes to the original M2MT-Net to ensure a fair comparison. As indicated in Table III, substituting the M2MT with a vanilla Transformer deteriorates the performance by 0.56 dB to 29.29 dB. Opting for a convolution results in a further decline, with a drop of 0.83 dB to 29.02 dB. These outcomes affirm M2MT’s efficacy as a feature extractor in the spatial subspace compared to other alternatives when paired with its angular subspace counterpart.
We also train a M2MT-Net variant with the angular Transformer replaced with a convolution. Surprisingly, this variant achieves 29.42 dB PSNR, which is comparable to the best competitor, LF-DET, and outperforms other Transformer-based competitors like LFT and EPIT. This underscores the robust adaptability of the M2MT, even when paired with less potent components in the angular subspace.
IV-G2 Correlation Tensor Channels
@@
TABLE IV: Ablation study of M2MT-Net’s correlation tensor channel number $C_{Cor}$. The best PSNR/SSIM are in bold.
Method FLOPs (G) Time (s) #Params. (M) Memory (GB) PSNR/SSIM
$C_{Cor}=64$ 31.21 2.04 2.460 5.04 29.60/0.9256
$C_{Cor}=96$ 32.50 2.05 3.198 5.08 29.67/0.9265
$C_{Cor}=128$ 33.85 2.09 3.986 5.10 29.85/0.9284
$C_{Cor}=160$ 35.24 2.11 4.824 5.12 29.72/0.9265
We evaluate the impact of varying the correlation tensor channel number $C_{Cor}$ on the performance and efficiency of M2MT-Net. Variants with $C_{Cor}\in\{64,96,160\}$ are trained and compared to the baseline $C_{Cor}=128$, with results detailed in Table IV.
The PSNR results indicate an increase with the increment of $C_{Cor}$ from 64 to 128, starting from 29.60 dB and peaking at 29.85 dB. However, further increasing $C_{Cor}$ to 160 leads to a slight decrease in PSNR to 29.72 dB, which may be attributed to overfitting and redundancy in the correlation tensor. This suggests that the optimal value for $C_{Cor}$ is 128.
Additionally, the impact of $C_{Cor}$ on the model’s memory and computational complexity is manageable. Increasing $C_{Cor}$ by 2.5 times from 64 to 160 results in a proportional 96.10% increase in the parameter number, yet the increments in FLOPs, inference time, and peak GPU memory usage are relatively modest at 12.91%, 3.43%, and 1.59%, respectively. These findings indicate that the correlation encoding and decoding processes are unlikely to be bottlenecks in the model’s efficiency.
V Conclusion and Future Work
@@
In this paper, we have revealed the prevalent challenge of subspace isolation caused by the One-to-One scheme and present the novel concept of Many-to-Many Transformers (M2MT) as a new scheme to address this issue. The proposed M2MT is empowered with complete access to all pixels across all SAIs in a LF image to capture comprehensive long-range correlation dependencies. With M2MT as a pivotal component, we have proposed a simple yet effective M2MT-Net for LFSR. Extensive experiments on various public datasets have demonstrated that M2MT-Net surpasses state-of-the-art methods in terms of reconstructed image quality while maintaining favorable computational and memory efficiency, making it a viable model for practical LF applications. Further, our analysis of angular consistency through LF depth estimation shows that M2MT-Net not only reconstructs finer visual details but also preserves and enhances the parallax structure of LF images. Its superiority is evidenced by visual interpretability in our in-depth analysis using the LAM technique, which highlights that M2MT involves a substantially broader range of pixels across wider SAIs beyond subspace isolation, signifying its truly global context and a more comprehensive modeling of correlation dependencies.
Looking ahead, there are some promising directions for improving M2MT in future works:
  1. 1.
    More Subspaces. Enhancing M2MT’s capacity by extending it to subspaces like the EPI can address large disparities, as seen in datasets like STFgantry [75], akin to EPIT [41]. Additionally, applying M2MT to the angular subspace could create a symmetric structure with its existing spatial counterpart. Nonetheless, two main challenges are anticipated: Firstly, the unique characteristics of these subspaces may require specific modifications to the Many-to-Many mechanism. Secondly, increasing model complexity becomes a significant concern when either spatial dimension, $W$ or $H$, is engaged as embeddings in the correlation encoding and decoding processes. Currently, these processes are achieved by manageable linear layers between $UVC$ and $C_{Cor}$. As $W\gg U$ and $H\gg V$, directly replacing $U$ or $V$ with $W$ or $H$ leads to unmanageable memory and computational complexities.
  2. 2.
    Light Field View Synthesis. M2MT holds potential for application in LF view synthesis. Particularly, some LF view synthesis methods [24, 29, 27] employ EPI-based strategies by extracting correlation features from EPIs or super-resolving EPIs to super-resolve the angular resolution. As discussed in the first point, enabling M2MT to operate within the EPI subspaces could effectively leverage its capabilities for this task.
  3. 3.
    Unified M2MT. It will be a compelling advancement to unify the spatial and angular Transformers into a single and holistic M2MT component. This integration would enable simultaneous and cohesive execution of spatial and angular self-attention processes, likely improving the model’s efficiency and effectiveness with compact spatial-angular feature extraction.